# Group-Relative REINFORCE Is Secretly an Off-Policy Algorithm: Demystifying Some Myths About GRPO and Its Friends

### v1

@@ Abstract
Off-policy reinforcement learning (RL) for large language models (LLMs) is attracting growing interest, driven by practical constraints in real-world applications, the complexity of LLM-RL infrastructure, and the need for further innovations of RL methodologies. While classic REINFORCE and its modern variants like Group Relative Policy Optimization (GRPO) are typically regarded as on-policy algorithms with limited tolerance of off-policyness, we present in this work a first-principles derivation for group-relative REINFORCE without assuming a specific training data distribution, showing that it admits a native off-policy interpretation. This perspective yields two general principles for adapting REINFORCE to off-policy settings: regularizing policy updates, and actively shaping the data distribution. Our analysis demystifies some myths about the roles of importance sampling and clipping in GRPO, unifies and reinterprets two recent algorithms — Online Policy Mirror Descent (OPMD) and Asymmetric REINFORCE (AsymRE) — as regularized forms of the REINFORCE loss, and offers theoretical justification for seemingly heuristic data-weighting strategies. Our findings lead to actionable insights that are validated with extensive empirical studies, and open up new opportunities for principled algorithm design in off-policy RL for LLMs. Source code for this work is available at https://github.com/modelscope/Trinity-RFT/tree/main/examples/rec_gsm8k.
11footnotetext: Equal contribution. Contact: chaorui@ucla.edu, chenyanxi.cyx@alibaba-inc.com22footnotetext: UCLA. Work done during an internship at Alibaba Group.33footnotetext: Alibaba Group.
@@ 1 Introduction
Alongside these developments, off-policy RL has been attracting growing interest. In the “era of experience” (Silver and Sutton, 2025), LLM-powered agents need to be continually updated through interaction with the environment. Practical constraints in real-world deployment and the complexity of LLM-RL infrastructure often render on-policy training impractical (Noukhovitch et al., 2025): rollout generation and model training can proceed at mismatched speeds, data might be collected from different policies, reward feedback might be irregular or delayed, and the environment may be too costly or unstable to query for fresh trajectories. Moreover, in pursuit of higher sample efficiency and model performance, it is desirable to go beyond the standard paradigm of independent rollout sampling, e.g., via replaying past experiences (Schaul et al., 2016; Rolnick et al., 2019; An et al., 2025), synthesizing higher-quality experiences based on auxiliary information (Da et al., 2025; Liang et al., 2025; Guo et al., 2025), or incorporating expert demonstrations into online RL (Yan et al., 2025; Zhang et al., 2025c) — all of which incur off-policyness.
However, the prominent algorithms in LLM-RL — Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024) — are essentially on-policy methods: as modern variants of REINFORCE (Williams, 1992), their fundamental rationale is to produce unbiased estimates of the policy gradient, which requires fresh data sampled from the current policy. PPO and GRPO can handle a limited degree of off-policyness via importance sampling, but require that the current policy remains sufficiently close to the behavior policy. Truly off-policy LLM-RL often demands ad-hoc analysis and algorithm design; worse still, as existing RL infrastructure (Sheng et al., 2024; Hu et al., 2024; von Werra et al., 2020; Wang et al., 2025; Pan et al., 2025; Fu et al., 2025a) is typically optimized for REINFORCE-style algorithms, their support for specialized off-policy RL algorithms could be limited. All these have motivated our investigation into principled and infrastructure-friendly algorithm design for off-policy RL.
@@
$=\frac{1}{K}\sum_{1\leq i\leq K}\sum_{1\leq t\leq|y_{i}|}(r_{i}-\overline{r})\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}^{t}\,|\,x,y_{i}^{<t})\quad\text{(token-wise)}.$ (1b)
@@ 2.1 Recap: on-policy interpretation via policy gradient theory
$\widetilde{g}(\bm{\theta};x)$ $={\mathbb{E}}_{y\sim\pi_{\textsf{b}}(\cdot|x)}\!\bigg[\big(r(x,y)-b(x)\big)\,\sum_{1\leq t\leq|y|}\frac{\pi_{\bm{\theta}}(y^{t}\,|\,x,y^{<t})}{\pi_{\textsf{b}}(y^{t}\,|\,x,y^{<t})}\,\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y^{t}\,|\,x,y^{<t})\bigg].$
Although this becomes a biased approximation of $\nabla_{\bm{\theta}}J(\bm{\theta};x)$, classical RL theory still offers policy improvement guarantees if $\pi_{\bm{\theta}}$ is sufficiently close to $\pi_{\textsf{b}}$ (Kakade and Langford, 2002; Fragkiadaki, 2018; Schulman et al., 2015, 2017; Achiam et al., 2017).
2.2 A new interpretation: REINFORCE is inherently off-policy
@@
where $\tau$ is a regularization coefficient. It is a well-known fact that the optimal policy $\pi$ for this surrogate objective satisfies the following (Nachum et al., 2017; Korbak et al., 2022; Rafailov et al., 2023; Richemond et al., 2024; Kimi-Team, 2025b): for any prompt $x$ and response $y$,
@@
Given a prompt $x$ and $K$ responses $y_{1},\dots,y_{K}$, we define the following mean-squared surrogate loss that enforces the consistency condition:
@@
Here, we normalize $a_{i}-a_{j}$ by $1+\tau$ to account for the loss scale. In theory, if this surrogate loss is defined by infinite samples with sufficient coverage of the action space, then its minimizer is the same as the optimal policy for the surrogate objective in Eq. (3).
@@
Figure 1 visualizes the proposed interpretation of what REINFORCE is actually doing. The curve going through $\bm{\theta}_{t}\rightarrow\bm{\theta}_{t+1}\rightarrow\widetilde{\bm{\theta}}_{t+1}\rightarrow\bm{\theta}^{\star}$ stands for the ideal optimization trajectory from $\bm{\theta}_{t}$ to the optimal policy model parametrized by $\bm{\theta}^{\star}$, if the algorithm solves each intermediate surrogate objective $J(\bm{\theta};\pi_{\bm{\theta}_{t}})$ / surrogate loss $\widehat{L}(\bm{\theta};\pi_{\bm{\theta}_{t}})$ exactly at each iteration $t$. In comparison, REINFORCE is effectively taking a single gradient step of the surrogate loss and immediately moving on to the next iteration $\bm{\theta}_{t+1}$ with a new surrogate objective.
@@
Figure 1: A visualization of our off-policy interpretation for group-relative REINFORCE. Here $\widehat{L}(\bm{\theta};\pi_{\bm{\theta}_{t}})={\mathbb{E}}_{x\sim\widehat{D}}[\widehat{L}(\bm{\theta};x,\pi_{\bm{\theta}_{t}})]$, where $\widehat{D}$ is the sampling distribution for prompts, and $\widehat{L}(\bm{\theta};x,\pi_{\bm{\theta}_{t}})$ is the surrogate loss defined in Eq. (6) for a specific prompt $x$.
@@ 4 Rethinking the rationales behind recent RL algorithms
F1 GRPO’s effectiveness in off-policy settings stems from clipping as regularization rather than importance sampling. A wider clipping range than usual often accelerates training without harming stability. Section 4.1, Figures 3, 4, 6, 9, 10
F2 Kimi’s OPMD and Meta’s AsymRE can be interpreted as REINFORCE loss + regularization loss, a perspective that is complementary to the rationales in their original papers. Section 4.2, Figure 11
F3 Data-oriented heuristics — such as dropping excess negatives or up-weighting high-reward rollouts — fit naturally into our off-policy view and show strong empirical performance. Section 4.3, Figures 5, 6, 7
@@
We primarily consider two off-policy settings in our experiments, specified by the sync_interval and sync_offset parameters in the Trinity-RFT framework (Pan et al., 2025). sync_interval specifies the number of generated rollout batches (each corresponding to one gradient step) between two model synchronization operations, while sync_offset specifies the relative lag between the generation and consumption of each batch. These parameters can be deliberately set to large values in practice, for improving training efficiency via pipeline parallelism and reduced frequency of model synchronization. In addition, $\texttt{sync\_offset}>1$ serves to simulate realistic scenarios where environmental feedback could be delayed. We also consider a stress-test setting that only allows access to offline data generated by the initial policy model. See Figure 2 for an illustration of these off-policy settings, and Appendix B.2 for further details.
We conduct experiments on math reasoning tasks like GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021) and Guru (math subset) (Cheng et al., 2025), as well as tool-use tasks like ToolACE (Liu et al., 2025a). We consider models of different families and scales, including Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct (Qwen-Team, 2025), Llama-3.1-8B-Instruct, and Llama-3.2-3B-Instruct (Dubey et al., 2024). Additional experiment details can be found in Appendix B.
@@ 4.1 Demystifying myths about GRPO
Figure 3: Empirical results for REC algorithms on GSM8k with Qwen2.5-1.5B-Instruct. Training reward curves are smoothed with a running-average window of size 3. Numbers in the legend denote clipping parameters $\epsilon_{\textsf{low}},\epsilon_{\textsf{high}}$.
@@
Figure 3 presents GSM8k results with Qwen2.5-1.5B-Instruct in various off-policy settings. REC-OneSide-IS/NoIS and GRPO (with the same $\epsilon_{\textsf{low}}=\epsilon_{\textsf{high}}=0.2$) have nearly identical performance, indicating that importance sampling is non-essential, whereas the collapse of REINFORCE highlights the critical role of clipping. Radically enlarging $(\epsilon_{\textsf{low}},\epsilon_{\textsf{high}})$ to $(0.6,2.0)$ accelerates REC-OneSide-NoIS without compromising stability in both sync_interval = 20 and sync_offset = 10 settings. Similar patterns also appear in Figure 4 (ToolAce with Llama-3.2-3B-Instruct) and other results in Appendix B. As for the stress-test (“offline”) setting, Figure 3 reveals an intrinsic trade-off between the speed and stability of policy improvement, motivating future work toward better algorithms that achieve both.
@@ 4.2 Understanding Kimi’s OPMD and Meta’s AsymRE
$\widehat{L}\big(\bm{\theta};x,\{y_{i},r_{i}\}_{1\leq i\leq K}\big)$ $=-\frac{1}{K}\sum_{i\in[K]}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}\,|\,x)+\tau\cdot R\big(\bm{\theta};x,\{y_{i},r_{i}\}_{1\leq i\leq K}\big),$
@@
With $K$ responses generated by $\pi_{\textsf{old}}=\pi_{\bm{\theta}_{t}}$, the term $\tau\log Z(x,\pi_{\bm{\theta}_{t}})$ can be approximated by a finite-sample estimate $\tau\log(\sum_{i}e^{r_{i}/\tau}/K)$, which can be further approximated by the mean reward $\overline{r}=\sum_{i}r_{i}/K$ if $\tau$ is large. With these approximations, the gradient of $\widetilde{L}$ becomes equivalent to that of the following loss (which is the final version of Kimi’s OPMD):
$\widehat{L}=-\frac{1}{K}\sum_{1\leq i\leq K}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}\,|\,x)+\frac{\tau}{2K}\sum_{1\leq i\leq K}\Big(\log\pi_{\bm{\theta}}(y_{i}\,|\,x)-\log\pi_{\textsf{old}}(y_{i}\,|\,x)\Big)^{2}.$
@@
AsymRE (Arnal et al., 2025) modifies REINFORCE by tuning down the baseline (from $\overline{r}$ to $\overline{r}-\tau$) in advantage calculation, which was motivated by the intuition of prioritizing learning from positive samples and justified by multi-arm bandit analysis in the original paper. We offer an alternative interpretation for AsymRE by rewriting its loss function:
$\widehat{L}$ $=-\frac{1}{K}\sum_{i}\Big(r_{i}-(\overline{r}-\tau)\Big)\log\pi_{\bm{\theta}}(y_{i}\,|\,x)=-\frac{1}{K}\sum_{i}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}\,|\,x)-\frac{\tau}{K}\sum_{i}\log\pi_{\bm{\theta}}(y_{i}\,|\,x).$

### v2

# arXiv:2509.24203v2[cs.LG] 01 Mar 2026

@@ Abstract
Off-policy reinforcement learning (RL) for large language models (LLMs) is attracting growing interest, driven by practical constraints in real-world applications, the complexity of LLM-RL infrastructure, and the need for further innovations of RL methodologies. While classic REINFORCE and its modern variants like Group Relative Policy Optimization (GRPO) are typically regarded as on-policy algorithms with limited tolerance of off-policyness, we present in this work a first-principles derivation for group-relative REINFORCE — a REINFORCE variant that uses the within-group mean reward as the baseline for advantage calculation — without assuming a specific training data distribution, showing that it admits a native off-policy interpretation. This perspective yields two general principles for adapting REINFORCE to truly off-policy settings: regularizing policy updates, and actively shaping the data distribution. Our analysis demystifies some myths about the roles of importance sampling and clipping in GRPO, unifies and reinterprets two recent algorithms — Online Policy Mirror Descent and Asymmetric REINFORCE — as regularized forms of the REINFORCE loss, and offers theoretical justification for seemingly heuristic data-weighting strategies. Our findings lead to actionable insights that are validated with extensive empirical studies, and open up new opportunities for principled algorithm design in off-policy RL for LLMs. Source code for this work is available at https://github.com/agentscope-ai/Trinity-RFT/tree/main/examples/rec_gsm8k.
@@ 1 Introduction
Alongside these developments, off-policy RL has been attracting growing interest. In the “era of experience” (Silver & Sutton, 2025), LLM-powered agents need to be continually updated through interaction with the environment. Practical constraints in real-world deployment and the complexity of LLM-RL infrastructure often render on-policy training impractical (Noukhovitch et al., 2025): rollout generation and model training can proceed at mismatched speeds, data might be collected from different policies, reward feedback might be irregular or delayed, and the environment may be too costly or unstable to query for fresh trajectories. Moreover, in pursuit of higher sample efficiency and model performance, it is desirable to go beyond the standard paradigm of independent rollout sampling, e.g., via replaying past experiences (Schaul et al., 2016; Rolnick et al., 2019; An et al., 2025), synthesizing higher-quality experiences based on auxiliary information (Da et al., 2025; Liang et al., 2025; Guo et al., 2025), or incorporating expert demonstrations into online RL (Yan et al., 2025; Zhang et al., 2025c) — all of which incur off-policyness.
However, the prominent algorithms in LLM-RL — Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024) — are essentially on-policy methods: as modern variants of REINFORCE (Williams, 1992), their fundamental rationale is to produce unbiased estimates of the policy gradient, which requires fresh data sampled from the current policy. PPO and GRPO can handle a limited degree of off-policyness via importance sampling, but require that the current policy remains sufficiently close to the behavior policy. Truly off-policy LLM-RL often demands ad-hoc analysis and algorithm design; worse still, as existing RL infrastructure (Sheng et al., 2024; Hu et al., 2024; von Werra et al., 2020; Wang et al., 2025b; Pan et al., 2025; Fu et al., 2025a) is typically optimized for REINFORCE-style algorithms, their support for specialized off-policy RL algorithms could be limited. All these have motivated our investigation into principled and infrastructure-friendly algorithm design for off-policy RL.
@@
$=\frac{1}{K}\sum_{1\leq i\leq K}\sum_{1\leq t\leq|y_{i}|}(r_{i}-\overline{r})\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}^{t}\,|\,x,y_{i}^{<t})\quad\text{(token-wise)}$ (1b)
@@ 2.1 Recap: on-policy interpretation via policy gradient theory
$\widetilde{g}(\bm{\theta};x)$ $={\mathbb{E}}_{y\sim\pi_{\textsf{b}}(\cdot|x)}\!\bigg[\big(r(x,y)-b(x)\big)\,\sum_{1\leq t\leq|y|}\frac{\pi_{\bm{\theta}}(y^{t}\,|\,x,y^{<t})}{\pi_{\textsf{b}}(y^{t}\,|\,x,y^{<t})}\,\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y^{t}\,|\,x,y^{<t})\bigg]$
Although this becomes a biased approximation of $\nabla_{\bm{\theta}}J(\bm{\theta};x)$, classical RL theory still offers policy improvement guarantees if $\pi_{\bm{\theta}}$ is sufficiently close to $\pi_{\textsf{b}}$ (Kakade & Langford, 2002; Fragkiadaki, 2018; Schulman et al., 2015; 2017; Achiam et al., 2017).
2.2 A new off-policy interpretation for group-relative REINFORCE
@@
where $\tau$ is a regularization coefficient. It is a well-known fact that the optimal policy $\pi$ for this surrogate objective satisfies the following (Ziebart et al., 2008) (Nachum et al., 2017; Korbak et al., 2022; Rafailov et al., 2023; Richemond et al., 2024; Kimi-Team, 2025b): for any prompt $x$ and response $y$,
@@
Given a prompt $x$ and $K$ responses $y_{1},\dots,y_{K}$, we define the following mean-squared surrogate loss that enforces the consistency condition, as done in prior works (Gao et al., 2024; Flet-Berliac et al., 2024):
@@
Here, we normalize $a_{i}-a_{j}$ by $1+\tau$ to account for the loss scale. In theory, if this surrogate loss is defined by infinite samples with sufficient coverage of the action space (Song et al., 2024), then its unique minimizer is the same as the optimal policy for the surrogate objective in Eq. (3).
@@
Figure 1 visualizes the proposed interpretation of what REINFORCE is actually doing. The curve going through $\bm{\theta}_{t}\rightarrow\bm{\theta}_{t+1}\rightarrow\widetilde{\bm{\theta}}_{t+1}\rightarrow\bm{\theta}^{\star}$ stands for the ideal optimization trajectory from $\bm{\theta}_{t}$ to the optimal policy model $\bm{\theta}^{\star}$, if the algorithm solves each intermediate surrogate objective $J(\bm{\theta};\pi_{\bm{\theta}_{t}})$ / surrogate loss $\widehat{L}(\bm{\theta};\pi_{\bm{\theta}_{t}})$ exactly at each iteration $t$. In comparison, REINFORCE is effectively taking a single gradient step of the surrogate loss and immediately moving on to the next iteration $\bm{\theta}_{t+1}$ with a new surrogate objective.
@@
Figure 1: A visualization of our off-policy interpretation for group-relative REINFORCE. Here $\widehat{L}(\bm{\theta};\pi_{\bm{\theta}_{t}})={\mathbb{E}}_{x\sim\widehat{D}}[\widehat{L}(\bm{\theta};x,\pi_{\bm{\theta}_{t}})]$, where $\widehat{D}$ is the sampling distribution for prompts and $\widehat{L}(\bm{\theta};x,\pi_{\bm{\theta}_{t}})$ is the loss defined in Eq. (6).
@@ 4 Rethinking the rationales behind recent RL algorithms
F1 GRPO’s effectiveness in off-policy settings stems from clipping as regularization rather than importance sampling. A wider clipping range than usual often accelerates training without harming stability. Section 4.1, Figures 2, 3, 5, 8, 9
F2 Kimi’s OPMD and Meta’s AsymRE can be interpreted as REINFORCE loss + regularization loss, complementary to the rationales in their original papers. Section 4.2, Figure 10
F3 Data-oriented heuristics — such as dropping excess negatives or up-weighting high-reward rollouts — fit naturally into our off-policy view and show strong empirical performance. Section 4.3, Figures 4, 5, 11
@@
We conduct experiments with the Trinity-RFT framework (Pan et al., 2025), and control off-policyness with the sync_interval (frequency of model synchronization) and sync_offset (lag between rollout generation and training) parameters. Larger values of these parameters improve efficiency (via pipeline parallelism) at the cost of off-policyness; in addition, $\texttt{sync\_offset}>0$ simulates delayed environmental feedback in practical scenarios. We also include a stress-test setting that only allows access to offline data generated by the initial policy model. Our experiments cover math reasoning tasks like GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), Guru-Math (Cheng et al., 2025), and tool-use tasks like ToolACE (Liu et al., 2025a). LLMs under consideration include Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct (Qwen-Team, 2025a), Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct (Dubey et al., 2024), and Qwen3-30B-A3B (Qwen-Team, 2025b). Further details can be found in Appendix B.
@@ 4.1 Demystifying myths about GRPO
Figure 2: Empirical results for REC algorithms on GSM8k with Qwen2.5-1.5B-Instruct. Training reward curves are smoothed with a running-average window of size 3. Numbers in the legend denote clipping parameters $\epsilon_{\textsf{low}},\epsilon_{\textsf{high}}$. The “mixed” setting adopts sync_interval = 16 and sync_offset = 8.
@@
We conduct experiments to validate Finding F1 regarding the roles of clipping (with a small or large clipping range) and importance sampling in GRPO. Figure 2 presents GSM8k results with Qwen2.5-1.5B-Instruct in various off-policy settings. REC-OneSide-IS/NoIS and GRPO (with the same $\epsilon_{\textsf{low}}=\epsilon_{\textsf{high}}=0.2$) have nearly identical performance, indicating that importance sampling is non-essential, whereas the collapse of REINFORCE highlights the critical role of clipping. Radically enlarging $(\epsilon_{\textsf{low}},\epsilon_{\textsf{high}})$ to $(0.6,2.0)$ accelerates REC-OneSide-NoIS without compromising stability in all considered settings (except “offline”). Similar patterns also appear in Figure 3 (ToolAce with Llama-3.2-3B-Instruct) and other results in Appendix B. As for the stress-test (“offline”) setting, Figure 2 reveals an intrinsic trade-off between learning speed and stability, motivating future work toward better algorithms that achieve both. We hypothesize that sequence-level importance sampling in GSPO (Zheng et al., 2025) could be non-essential as well; interested readers may refer to our preliminary results in Appendix B.7 that support this prediction.
@@ 4.2 Understanding Kimi’s OPMD and Meta’s AsymRE
$\widehat{L}\big(\bm{\theta};x,\{y_{i},r_{i}\}_{1\leq i\leq K}\big)$ $=-\frac{1}{K}\sum_{i\in[K]}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}\,|\,x)+{\beta}\cdot R\big(\bm{\theta};x,\{y_{i},r_{i}\}_{1\leq i\leq K}\big),$
@@
With $K$ responses generated by $\pi_{\textsf{old}}=\pi_{\bm{\theta}_{t}}$, the term $\tau\log Z(x,\pi_{\bm{\theta}_{t}})$ can be approximated by a finite-sample estimate $\tau\log(\sum_{i}e^{r_{i}/\tau}/K)$ (Brantley et al., 2025), which can be further approximated by the mean reward $\overline{r}=\sum_{i}r_{i}/K$ if $\tau$ is large. With these approximations, the gradient of $\widetilde{L}$ becomes equivalent to that of the following loss (which is the final version of Kimi’s OPMD):
$\widehat{L}=-\frac{1}{K}\sum_{1\leq i\leq K}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}\,|\,x)+\frac{{\beta}}{2K}\sum_{1\leq i\leq K}\Big(\log\pi_{\bm{\theta}}(y_{i}\,|\,x)-\log\pi_{\textsf{old}}(y_{i}\,|\,x)\Big)^{2},\,\,{\text{where}}\,\,{\beta=\tau}.$
@@
AsymRE (Arnal et al., 2025) modifies REINFORCE by tuning down the baseline (from $\overline{r}$ to $\overline{r}-{\beta}$) in advantage calculation, which was motivated by the intuition of prioritizing learning from positive samples and justified by multi-arm bandit analysis in the original paper. We offer an alternative interpretation for AsymRE by rewriting its loss function:
$\widehat{L}$ $=-\frac{1}{K}\sum_{i}\Big(r_{i}-(\overline{r}-{\beta})\Big)\log\pi_{\bm{\theta}}(y_{i}\,|\,x)=-\frac{1}{K}\sum_{i}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}\,|\,x)-\frac{{\beta}}{K}\sum_{i}\log\pi_{\bm{\theta}}(y_{i}\,|\,x).$